OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

Slope intercept form: \[y - y _{1} = m ( x- x _{1})\] Plug in 4 for slope and (-3,-1) as the points, where -3 is y1 and -1 is x2

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

For an example, if i had a slope of 3, and the points i had were (-4,-2) This is how it would look: y - -2 = 3(x - - 4) y + 2 = 3(x+4) Do you get it?

OpenStudy (anonymous):

just put x=-3 and y=-1 in your equations

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so y-1= 4 (x- - 3)

OpenStudy (anonymous):

y+1 = 4 (x+3)

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

so its y= 4x+3

OpenStudy (anonymous):

kelbbb